### v1

# The Relative Contributions of Accretion Disk versus Jet to the Optical and Mid-infrared Variability of Seyfert Galaxies

Abstract
We performed a comprehensive analysis of flux and color variability in a redshift-matched sample of Seyfert galaxies, comprising 23 gamma-ray-detected narrow-line Seyfert 1 galaxies (gNLS1s), 190 non-gamma-ray-detected narrow-line Seyfert 1 galaxies (ngNLS1s), and 10 gamma-ray-detected broad-line Seyfert 1 galaxies (gBLS1s). Utilizing multi-band light curves from the Zwicky Transient Facility (ZTF) in g, r, and i bands, along with mid-infrared (MIR) observations in W1 and W2 bands from the Wide-field Infrared Survey Explorer (WISE), we observed that gBLS1s exhibit more significant variability than gNLS1s, while ngNLS1s display minimal variability across both optical and MIR wavelengths. The pronounced variability in gBLS1s may be attributed to a more closely aligned jet relative to the observer’s line of sight or their comparatively lower accretion rates. In contrast, the subdued variability in ngNLS1s suggests that their flux changes are primarily driven by accretion disk instabilities. A significant correlation between optical and MIR variability amplitudes found here supports the reprocessing scenario, wherein variations in the accretion disk emission are re-emitted by surrounding dust. Furthermore, our long-term color variability analysis revealed a stronger redder-when-brighter (RWB) trend in approximately 35%, 61%, and 80% of gNLS1s, ngNLS1s, and gBLS1s, respectively, in optical wavelength, strengthens the reprocessing scenario with the observed trend of stronger RWB for approximately 68%, 96%, and 30%, respectively, in MIR wavelength. The prevalent RWB trend observed in both optical and MIR wavelengths from the current sample of Seyfert galaxies on the longer time scales is likely associated with accretion disk instabilities.
keywords:
surveys – galaxies: active – galaxies: jets – -ray-galaxies: photometry – galaxies: Seyfert – gamma-rays: galaxies1 Introduction
Active galactic nuclei (AGNs) are among the most energetic extragalactic sources, with bolometric luminosities reaching up to 1048 erg s-1 (Woo & Urry, 2002). These extraordinary objects are powered by the accretion of gas onto supermassive black holes (SMBHs) at the centers of their host galaxies, with SMBH masses ranging from 106 to 1010 M⊙. This accretion process is the primary mechanism driving their immense energy output, making AGNs the most luminous objects in the universe (Lynden-Bell, 1969; Rees, 1984; Koratkar & Blaes, 1999; Bischetti et al., 2017). AGN models propose that, along with the central SMBH, an accretion disk and a surrounding dusty torus are present (Antonucci, 1993; Urry & Padovani, 1995). This structure is supported by characteristic features observed in the broadband spectral energy distribution (SED), including the big blue bump (BBB) and the mid-infrared (MIR) bump. The BBB is believed to originate from thermal emission produced by the accretion disk (Shakura & Sunyaev, 1973), while the mid-IR bump is attributed to emission from the dusty torus, as revealed by dust reverberation mapping studies (Suganuma et al., 2006; Koshida et al., 2014; Mandal et al., 2018).
Among the defining characteristics of AGNs, flux variability across the electromagnetic spectrum stands out as one of the most significant traits (Wagner & Witzel, 1995; Ulrich et al., 1997). Such variability, observed on timescales ranging from hours to years across wavelengths from radio to -ray, has been extensively utilized to investigate emitting regions, the intricate nature of the central engine, and accretion processes, phenomena (see, Urry & Padovani, 1995; Wagner & Witzel, 1995; Ulrich et al., 1997; Czerny et al., 2008; Rakshit et al., 2015). Among these, optical variability has been a main focus of study since the discovery of AGNs, as it provides crucial insights into the physics of accretion disks and jets (Wagner & Witzel, 1995; Ghosh et al., 2000; Villata et al., 2004; Gupta et al., 2008; Rani et al., 2010; Gu & Ai, 2011; Ikejiri et al., 2011; Bonning et al., 2012; Mao & Zhang, 2016). In addition to optical studies, the investigation of infrared (IR) variability has proven equally important. Infrared echoes, which arise in response to variations in optical continuum, provide vital information about the morphology of the dust torus and the accretion mechanisms that power the AGNs (Maune et al., 2013). Moreover, IR wavelength presents unique advantages over optical wavelength, such as (1) the IR continuum being unaffected by the presence of strong emission lines in low-redshift AGNs, and (2) reduced sensitivity to dust extinction. Together, multiwavelength variability studies across the optical and IR bands provide a comprehensive understanding of the central regions and the processes shaping AGN emission. In particular, changes in the brightness of AGNs are often accompanied by spectral variations, which can be explored through color-magnitude or spectral index–magnitude correlations (e.g., see Ciprini et al., 2003). Examining spectral variability is particularly important, as it can help to pinpoint the origin of brightness fluctuations, even when the analysis is confined to the optical and IR bands (Vagnetti et al., 2003).
Optical variability studies on large samples of luminous AGNs in optical wavelength have resulted in various interesting relationships between the amplitude and timescale of variability with physical parameters of AGN such as black hole mass, Fe ii strength, accretion rate, bolometric luminosity, redshift, etc. (e.g., di Clemente et al., 1996; Vanden Berk et al., 2004; Kelly et al., 2009; MacLeod et al., 2010; Meusinger et al., 2011; Rakshit & Stalin, 2017; Li et al., 2018), providing new insights into the accretion process in AGNs. However, it has been limited to a few sources only for the low luminous part of AGNs such as narrow-line Seyfert 1 galaxies (NLS1s), characterized based upon the flux ratio of forbidden line with Balmer line (H) 3, and narrow width of Balmer emission line with full width at half maximum (FWHM) 2000 km s-1 (Osterbrock & Pogge, 1985; Goodrich et al., 1989). For instance, a few studies on radio-loud and radio-quiet NLS1s have been performed on the intra-night timescale (Paliya et al., 2013; Kshama et al., 2017; Ojha, 2022; Ojha et al., 2022, 2024) but very limited studies are available on longer time scales from days to year (Young et al., 1999; Miller et al., 2000; Klimek et al., 2004; Doroshenko et al., 2006; Maune et al., 2013). Similarly, scarce studies are available for NLS1s in the IR band (Jiang et al., 2012). On the other hand, NLS1s have a higher accretion rate, strong Fe ii emission, stronger X-ray variability, steeper X-ray spectra, lower black hole mass, and lower optical variability compared to their broad-line counterparts, namely broad-line Seyfert 1 galaxies (BLS1s; Leighly, 1999; Véron-Cetty et al., 2001; Grupe & Mathur, 2004; Zhou et al., 2006; Xu et al., 2012; Ojha et al., 2020). Despite the intrinsic differences between NLS1s and BLS1s, gamma-ray-detected Seyfert galaxies, including both NLS1s and BLS1s, are of particular interest because they exhibit optical variability on short timescales, similar to that observed in blazars (Ojha et al., 2022). This suggests a significant contribution from relativistic jets in these sources, raising the question whether the observed optical and MIR variability arises primarily from jet activity or from thermal processes in the accretion disk.
Although several studies have explored the variability of Seyfert galaxies, key gaps remain in our understanding of their multi-wavelength variability characteristics. For example, Rakshit & Stalin (2017) investigated long-term optical variability of the redshift-matched sample of NLS1s and BLS1s using Catalina Real-Time Transient Survey (CRTS) data, and Rakshit et al. (2019) investigated MIR flux and color variability in NLS1s using data from the Wide-Field Infrared Survey Explorer (WISE). However, a comprehensive and systematic investigation of flux and color variability from optical to MIR wavelengths, specifically comparing gamma-ray-detected NLS1s (gNLS1s), non-gamma-ray NLS1s (ngNLS1s), and gamma-ray-detected BLS1s (gBLS1s) is still lacking. Such a study is crucial for disentangling the relative contributions of jet and disk emissions in these systems, thereby advancing our understanding of the physical mechanisms driving variability across the electromagnetic spectrum. Given the availability of new catalog of NLS1s (Paliya et al., 2024), which is a factor of two times larger than the previous catalog given by Rakshit et al. (2017) and high cadence ( 1 day) photometric data, available in Zwicky Transient Facility (hereafter, ZTF; Bellm et al., 2019) for g, r, and i bands along-with MIR data from WISE (Wright et al., 2010), strongly motivate us to do a systematic study of flux and color variability for a sample of gNLS1, ngNLS1s, and gBLS1s across optical to MIR wavelengths. Therefore, for the optical study of gNLS1s, ngNLS1s, and gBLS1s, we used the ZTF (see, Bellm et al., 2019; Masci et al., 2019), which is an optical time-domain survey utilizing the 48-inch Schmidt Telescope at Palomar Observatory. It is equipped with a 6k6k mosaic CCD camera that provides a field of view (FoV) of 47 deg2 with a readout time of 8 sec (Graham et al., 2019) and scans the northern sky in g, r, and i bands with an average cadence up to 1 day.
Additionally, for the MIR study of the current samples, we utilized data from WISE, which surveys the entire sky in four MIR bands W1, W2, W3, and W4 centered at 3.4 m, 4.6 m, 12 m, and 22 m, respectively, with angular resolutions of 6.1, 6.4, 6.5, and 12 arcseconds, respectively. WISE has a FoV of 47 arcmin2 and completes a full-sky scan approximately every six months. During each visibility window, it continuously observes sources for at least one day, occasionally longer, with a cadence of 90 minutes. Observations were conducted during two primary phases: (i) the WISE Full Cryogenic and Near-Earth Object WISE (NEOWISE) Post-Cryo Missions (2010–2011), which covered three visibility windows in all four bands, with photometric data archived in the “AllWISE Multi-Epoch Photometry (MEP) Database”; and (ii) the NEOWISE Reactivation Mission (NEOWISE-R; Mainzer et al., 2014), spanning 2014–2024, which focused on the W1 and W2 bands across eight visibility windows, with data stored in the “NEOWISE-R single exposure (L1b) source table.
The format of this paper is as follows. In Sect. 2, we describe our sample selection and data compilation. Sect. 3 provides details of our methodologies used for the analysis of the data. The results of the current work, followed by discussion, are given in Sect. 4. The conclusions of the current work are described in Sect. 5.
2 Sample selection
For this study, we used the most recent catalog of NLS1 galaxies, comprising 22,656 sources (Paliya et al., 2024). To construct a sample of ngNLS1s, we first compiled all known gNLS1s identified by the Fermi Large Area Telescope (LAT) from the literature, including the 21 sources listed in Table 1 of Paliya et al. (2024). This compilation yielded a total of 25 gNLS1s. Applying a redshift cut of to ensure adequate signal-to-noise ratios (S/N) for variability analysis using data from the ZTF and WISE reduced the gNLS1 sample to 23 sources. In parallel, a control sample of 45 gBLS1s was selected from Paliya et al. (2024), also subject to the same redshift constraint of to maintain consistency in data quality across samples. Thus, finally, we identified 22,635 ngNLS1s, 23 gNLS1s, and 45 gBLS1s. Next, we prepared redshift-matched samples to enable a fair comparison. The resulting matched samples consist of 23 gNLS1s, 190 ngNLS1s, and 10 gBLS1s whose cumulative probability distribution for rest-frame redshifts is presented in Fig. 1 . Finally, a Kolmogorov–Smirnov (K-S) test performed on the redshift-matched samples yielded a null hypothesis probability () of approximately 21% for gNLS1s versus ngNLS1s and 28% for gNLS1s versus gBLS1s.
Out of different subsamples of Seyfert galaxies used in the current work, gNLS1s and gBLS1s can exhibit hour-like time-scale variability, which can create an obstacle in studying the true spectral behavior of these sources in the absence of simultaneous multi-band data, both in optical and MIR wavebands. Therefore, for the color variability study, we selected sources that have quasi-simultaneous observations within 30 minutes in different bands with at least five data points. To make it consistent, we have applied the same criterion of quasi-simultaneous observations to the sample of ngNLS1s. This has resulted in a sample of 17 gNLS1s, 128 ngNLS1s, and 10 gBLS1s for the g-r color study and 10 gNLS1s, 74 ngNLS1s, and 6 gBLS1s for the r-i color study in the optical wavelength. Similarly, samples of 22 gNLS1s, 176 ngNLS1s, and 10 gBLS1s resulted in for the W1-W2 color study in the mid-infrared wavelength.
2.1 Optical photometric data from ZTF
For the redshift-matched samples of 23 gNLS1s, 190 ngNLS1s, and 10 gBLS1s, we retrieved ZTF photometric data for the g, r, and i band in the 22nd ZTF public data release (see Masci et al., 2019) within a search radius of an arcsec to the target position from the NASA/IPAC IR Science Archive (IRSA111https://irsa.ipac.caltech.edu/Missions/ztf.html) using the application program interface (API). To eliminate any suspicious measurements that could mimic spurious variability, such as those arising from the independent calibration of ZTF light curves across different fields and CCD quadrants for the same filter (e.g., see van Roestel et al., 2021), we select only the light curve for a given filter corresponding to the observation ID with the maximum number of data points. Furthermore, to ensure the quality of the photometric data, we selected data points with catflags = 0, as recommended in the ZTF Science Data System Explanatory Supplement222https://irsa.ipac.caltech.edu/data/ZTF/docs/ztf explanatory supplement.pdf and the public data release notes333https://irsa.ipac.caltech.edu/data/ZTF/docs/releases/dr20/ztf release notes dr20.pdf, removed outliers by applying a 3 clipping on the entire data set and also excluded poor quality data points with uncertainties greater than 10% in magnitude. The precautions and filtering above enable us to determine small-magnitude variations in the generated g, r, and i band light curves. This has resulted in samples of 20 gNLS1s, 186 ngNLS1s, and 10 gBLS1s for the ZTF r-band. The total number of sources with available data in ZTF g, r, and i bands for different classes of Seyfert galaxies is tabulated in Table 1. The obtained lightcurves of Seyfert galaxies have magnitude data points at a cadence of as high as 1 day for sources in their field with a maximum time baseline of 2000 days. Representative light curves of the r-band for one of the gNLS1-J032441.28341045.1 (left), ngNLS1-J221541.64301435.7 (middle) and gBLS1- J095649.88251516.1 (right) from the current sample are shown in Fig. 2.
2.2 Mid-infrared data from WISE
For the redshift-matched samples of 23 gNLS1s, 190 ngNLS1s, and 10 gBLS1s, we retrieved MIR data from the MEP database and the NEOWISE-R single exposure (L1b) source database from the NASA/IPAC IR Science Archive444https://irsa.ipac.caltech.edu/Missions/wise.html, giving us data from 2010 to 2024 with a gap of about 3 years between 2011-2014 due to the depletion of the solid hydrogen cryogen used to cool the W3 and W4 instruments. Next, we limited our MIR study on the current sample to the W1 and W2 bands only due to the availability of the data for the W3 and W4 bands between 2010 and 2011, with many photometric points marked as ‘null’. Furthermore, both databases were screened to exclude bad photometric measurements, adopting the criteria (i) the reduced value of for the single exposure profile fit must be below 5 for both the W1 and W2 bands (that is, w1rchi2 5 and w2rchi2 5). (ii) The number of points spread function (PSF) components used in the profile fit for a source should be fewer than 3 (nb 3). (iii) Single-exposure images must have the highest quality (qifact = 1), be free from known artifacts (ccflags = ‘0000’), and not involve active d-eblending (na = 0). (iv) The number of photometric data points available in the W1 and W2 bands to be 5. (v) Also removed outliers by applying a 3 clipping on the whole data set.
With the above constraints, we arrived at final samples of 22 gNLS1s, 178 ngNLS1s, and 10 gBLS1s for their study of MIR flux variability. The total number of sources with available data in the WISE W1 and W2 bands for different classes of Seyfert galaxies is tabulated in Table 1. Furthermore, upon visually inspecting the generated light curves for the W1 and W2 bands, we observed that many light curves included photometric points separated by approximately 11 seconds in consecutive WISE epochs. Given that the WISE orbital period is roughly 1.5 hours, we averaged such photometric points with their nearest neighbors within this time frame to reconstruct the light curves for the entire sample. Representative light curves for one of the gNLS1-163915.80412833.7 (left), ngNLS1-J145758.08454119.2 (middle), and gBLS1-J103632.97220312.2 (right) from the current sample for the WISE W1 band (top panel) and W2 band (bottom panel) are shown in Fig. 3.
3 Methodology for analysis
Before proceeding further with the analysis, we first corrected the observed data of each source for the foreground galactic interstellar reddening and absorption, using the extinction correction (AV) values provided by NASA’s NED, following the method of (Schlafly & Finkbeiner, 2011). The extinction-corrected r-band light curves for a gNLS1, ngNLS1, and gBLS1 from the current sample are displayed in Fig. 2.
3.1 Optical flux variability
To confirm the presence/absence of variability in the final samples of gNLS1s, ngNLS1s, and gBLS1s (see Table 1) using the ZTF g, r, and i band light curves, we have used the standard -test (see, Goyal et al., 2012). A detailed explanation for this test is demonstrated in our previous papers (Ojha et al., 2022, 2024, and references therein).
In brief, following Goyal et al. (2012), -test can be written as
where is the variance and is the mean square (formal) rms errors of the ith data points in the light curves of the target Seyfert galaxy, and is the number of observations. A necessary requirement of this test is to use the correct rms errors on the photometric data points. However, in determining the instrumental magnitudes of the target in the ZTF data reduction pipeline, the Image Reduction and Analysis Facility (IRAF555http://iraf.noao.edu/) software package and the Dominion Astrophysical Observatory Photometry (DAOPHOT II666http://www.astro.wisc.edu/sirtf/daophot2.pdf) algorithm were used, which typically underestimate magnitude errors by a factor ranging from 1.3 to 1.75, as noted by previous studies (Sagar et al., 2004; Bachev et al., 2005). To correct for this underestimation, a value derived from 262 intra-night AGN monitoring sessions by Goyal et al. (2013) using a ground-based telescope of a meter class has been adopted. This correction is crucial for accurate photometric analysis, particularly for the rms errors of the photometric data points. In their study, Goyal et al. (2013) estimated the best fit value of to be 1.540.05, which accounts for the systematic underestimate of the magnitude errors by the data reduction software and ensures the reliability of the variability analysis. Therefore, in the current study, is considered due to the high precision of this estimate. To ensure a robust assessment of variability in the optical light curves of Seyfert galaxies, we employed the F-test, which compares the observed variance in the light curve to the expected variance due to measurement uncertainties, scaled by a factor in this work to use the rms errors of the photometric data points accurately. We adopted a stringent significance level of , corresponding to a confidence level of 99%, to minimize the probability of a Type I error (i.e., false identification of variability when none exists). In statistical terms, setting means that there is only a 1% chance of incorrectly rejecting the null hypothesis of no variability, thus ensuring that any detected variability is supported with high confidence. If a galaxy light curve is classified as variable at this threshold, we are 99% confident that the observed variability is intrinsic and not due to random fluctuations in the data points. The test statistic denoted as and defined in Eq. 1, was calculated for each light curve and compared with its critical F-value, , obtained from the F distribution for the appropriate degrees of freedom at the chosen significance level. The critical value depends on the degrees of freedom (here equal to the number of data points in the observed light curve) in the numerator and denominator. A galaxy is deemed variable (V) if its computed value of the F-test satisfies the condition , thus rejecting the null hypothesis and confirming the presence of statistically significant variability at the confidence level of 99%.
Furthermore, to estimate the level of brightness variations in the above samples, we have estimated (i) the peak-to-peak amplitude of flux variability () following the definition given by Heidt & Wagner (1996) and (ii) the fractional variability amplitude (Fvar) introduced by Vaughan et al. (2003) for all data sets.
Following the definition given by Heidt & Wagner (1996), the can be written as
where and are the maximum and minimum amplitude values in the light curve and is the mean square (formal) rms errors of individual data points scaled with the scaling factor , adopted from Goyal et al. (2013). Here, the precession of the light curve is considered for the error associated with .
Furthermore, to estimate the normalized intrinsic source variance free from the contribution expected from measurement errors, we follow Equation 3 introduced by Vaughan et al. (2003)
where, is the mean square error and given as follows:
is the sample variance for the flux measurements, . The associated error with can be computed following the given equation
For each gNLS1, ngNLS1, and gBLS1 galaxies, we have tabulated the duration of the light curve, number of data points in the light curve, variability status based on the -test, , and in Table 2.
3.1.1 Duty cycle of variability
AGNs do not necessarily exhibit flux variations during each epoch of their monitoring. Hence, it is more suitable to evaluate the duty cycle (DC) as the ratio of the time intervals when variability is observed to the total observing time rather than simply considering the fraction of variable objects. Following the definition proposed by Romero et al. (1999) and further utilized by Stalin et al. (2004) for intra-night variability, we computed the DC for the current sample of Seyfert galaxies using the following expression.
here, z represents the observed duration of the monitoring session after applying the redshift correction to the source. For the session, is assigned a value of 1 in Eq. 5 only when variability is observed, otherwise is set to 0.
In Table 1, we have tabulated the computed DC along with the mean and median values of and for the current sample of Seyfert galaxies across the g, r and i bands. It is important to note that when calculating the mean amplitude () and the median amplitude () for a sample, only light curves classified as variables were considered.
3.2 Mid-infrared flux variability
To quantitatively study the MIR variability of the current sample, we calculated the intrinsic amplitude of the variability () commonly used by several authors to study the MIR variability (e.g., see Rakshit et al., 2019; Anjum et al., 2020; Wang & Shi, 2020). The is the variance of the observed magnitudes after correcting for associated measurement errors, which can be calculated following Ai et al. (2010) and adopting the formalism used in Sesar et al. (2007).
where is the observed magnitude at the point, and is the weighted average. The can be expressed as:
where is calculated including systematic error () with the individual error () as follows:
The uncertainty associated with can be computed following the given equation
considering that the data points are independent and normally distributed, the uncertainty in can be approximated as:
here, systematic errors of 0.024 mags and 0.028 mags are used for the WISE W1 and W2 bands, respectively, as reported in Jarrett et al. (2011). Since the current sample has a redshift range between , therefore, spurious variability may occur if we do not correct for the redshift of the source, thus we calculated the rest frame by multiplying by . Here, we have considered a Seyfert galaxy to be variable if its computed value of is found to be 0.1. The for the individual source and its variability status are tabulated in Table 2 for the W1 and W2 bands and calculated values of , along with their associated uncertainties, for the samples of gNLS1s, ngNLS1s, and gBLS1s in the W1 and W2 bands are presented in Table 1.
3.3 Correlation between optical and mid-infrared variability amplitudes
It is expected that IR emission is reprocessed from optical/UV emission (Lu et al., 2016), meaning that the variability characteristics in the optical bands will be imprinted in the IR wavelength. To investigate such a scenario in the current sample, we selected Seyfert galaxies that exhibit optical variability and have a redshift-corrected variability amplitude in the MIR. For the r-band light curve, the criteria are satisfied by 14 gNLS1s, 8 ngNLS1s, and 9 gBLS1s in both the W1 and W2 bands. However, for the W2 band, only 6 ngNLS1s meet the criteria, while 8 ngNLS1s meet the criteria for the W1 band. Using this sample, we constructed correlation plots between and both and . A representative plot showing the correlated variability amplitudes of the r-band optical light curve with the MIR W1 and W2 bands for gNLS1s (left), ngNLS1s (middle), and gBLS1s (right) is presented in Fig. 4. To examine the relationship between MIR and optical variability amplitudes, we employed orthogonal distance regression (ODR) fitting on the data points in the variability amplitude correlation plot, accounting for uncertainties on both axes. Furthermore, we calculated the Pearson rank correlation coefficient () to quantify the correlation between the variability amplitudes for the gNLS1s, ngNLS1s, and gBLS1s samples across both the optical and MIR wavelengths. In this plot, the upper panels illustrate the correlation between and , while the lower panels show the correlation between and . For the r-band light curve, the results of the correlation analysis, such as slope, intercept, Pearson r for the W1 and W2 bands, are displayed in the upper left corner of each panel in Fig. 4.
For the optical wavelength | For the mid-infrared wavelength | |||||||||
Seyfert galaxies | Duty cycle | Mean | Median | Mean | Median | LC⋆ | Seyfert galaxies | Mean | Error | LC⋆ |
[no. of sources] | (in %) | () | () | Fvar () | Fvar | of | [no. of sources] | err_ | of | |
gNLS1 [22] | 68.80 | 1.350.12 | 1.000.10 | 1.110.03 | 0.830.02 | g-band | gNLS1 [22] | 0.287 | 0.059 | W1-band |
gNLS1 [20] | 82.67 | 1.250.10 | 0.950.08 | 1.010.02 | 0.850.01 | r-band | gNLS1 [22] | 0.324 | 0.070 | W2-band |
gNLS1 [16] | 65.61 | 0.840.08 | 0.630.07 | 0.900.04 | 0.750.04 | i-band | ngNLS1 [178] | 0.113 | 0.006 | W1-band |
ngNLS1 [171] | 6.42 | 0.860.11 | 0.740.10 | 0.560.07 | 0.530.04 | g-band | ngNLS1 [178] | 0.026 | 0.004 | W2-band |
ngNLS1 [186] | 5.18 | 0.750.10 | 0.660.09 | 0.420.05 | 0.440.03 | r-band | gBLS1 [10] | 0.447 | 0.091 | W1-band |
ngNLS1 [147] | 2.45 | 0.560.08 | 0.550.09 | 0.350.08 | 0.340.06 | i-band | gBLS1 [10] | 0.427 | 0.091 | W2-band |
gBLS1 [10] | 100 | 1.450.09 | 1.390.06 | 1.460.02 | 1.210.01 | g-band | —- | — | — | – |
gBLS1 [10] | 100 | 1.690.09 | 1.830.06 | 1.810.01 | 1.600.01 | r-band | —- | — | — | – |
gBLS1 [08] | 84.96 | 1.530.11 | 1.450.07 | 1.610.04 | 1.280.03 | i-band | —- | — | — | – |
†The mean and median values for the sample considering the light curves belonging to a variable (V) type only. ⋆Light curve | ||||||||||
The number of sources used to estimate the statistics are tabulated inside the bracket ‘[]’. |
Optical wavelength | Mid-infrared wavelength | ||||||||
SDSS Name | Redshift | Galaxy | Time | Optical | Variability | Variability | |||
(z) | type | (in days) | dpts† | status | status | ||||
For the g-band light curves | For the W1-band light curves | ||||||||
J000132.37211336.2 | 0.439 | gNLS1 | 514.781 | 120 | V | 0.7650.067 | 0.8670.022 | 0.1400.030 | V |
J000545.32063945.4 | 0.884 | ngNLS1 | 2207.987 | 381 | NV | 0.6600.121 | 0.4760.022 | 0.1540.016 | V |
J094420.44613550.1 | 0.791 | gBLS1 | 2044.211 | 1318 | V | 1.0430.084 | 0.8920.008 | 0.2550.038 | V |
For the r-band light curves | For the W2-band light curves | ||||||||
J000132.37211336.2 | 0.439 | gNLS1 | 999.672 | 146 | V | 1.0900.052 | 1.2840.016 | 0.3610.034 | V |
J000545.32063945.4 | 0.884 | ngNLS1 | 2198.960 | 506 | NV | 0.6730.116 | 0.4720.018 | 0.0660.047 | NV |
J094420.44613550.1 | 0.791 | gBLS1 | 2032.324 | 1475 | V | 0.9450.078 | 0.8040.007 | 0.1530.030 | V |
For the i-band light curves | |||||||||
J000132.37211336.2 | 0.439 | gNLS1 | 1239.762 | 47 | V | 0.8520.061 | 1.0780.033 | **** | **** |
J000545.32063945.4 | 0.884 | ngNLS1 | 1573.788 | 121 | NV | 0.6390.137 | 0.5330.044 | **** | **** |
J094420.44613550.1 | 0.791 | gBLS1 | 1211.898 | 147 | NV | 0.5540.104 | 0.3890.031 | **** | **** |
*Data is not available. †Number of data points in the optical band light curve. V: Variable; NV: Non-variable. | |||||||||
Notes: A portion of this table is presented here to display its form and content, however, the entire table is available in the online | |||||||||
electronic version. |
3.4 Color variability
The optical emission observed in AGNs arises from two primary components: the quasi-thermal emission originating from the accretion disk and the non-thermal synchrotron emission from the relativistic jet. Analyzing the color variability of the current sample can provide an effective way to differentiate these components that contribute to the overall flux. Similarly, IR emission, which is considered reprocessed optical/UV radiation (see, Storchi-Bergmann et al., 1992; Lu et al., 2016), offers a unique opportunity to examine the interplay of thermal and non-thermal processes in AGNs. However, to ensure accurate spectral analysis, simultaneous measurements in different bands must be obtained within a 30-minute interval, as described in Sect. 2. This criterion helps mitigate any uncertainties caused by temporal variations and ensures a reliable study of the spectral behavior. Consequently, the color variability analysis in both optical and MIR wavelengths is limited to a subset of sources that meet this stringent timing requirement. The details of the final samples satisfying these conditions are provided at the end of Sect. 2.
For the samples of gNLS1s, ngNLS1s, and gBLS1s mentioned in the last paragraph of Sect. 2, we generated color-magnitude plots, including g-r vs. r, r-i vs. i, and W1-W2 vs. W2, using the available quasi-simultaneous data across different optical and MIR bands. Here, the magnitudes of the longer wavelength are chosen on the X-axis, and the difference between the magnitudes of shorter and longer wavelengths is considered for the color to make the color-magnitude analysis consistent across optical and MIR wavelengths. Representative examples of these long-term best-fit color-magnitude plots for the gNLS1-J130522.74511640.2 (left), ngNLS1-J144019.82502201.7 (middle), and gBLS1-J110249.85525012.7 (right) are presented in Fig. 5, corresponding to g-r vs. r, r-i vs. i, and W1-W2 vs. W2, respectively, for top, middle, and bottom. To examine the relationship between color and magnitude, we employed ODR fitting on the data points in the color-magnitude plots, accounting for uncertainties in both axes. Furthermore, we calculated to quantify the correlation between color and magnitude for the gNLS1s, ngNLS1s, and gBLS1s samples across both optical and MIR wavelengths.
To qualitatively identify bluer-when-brighter (BWB) or redder-when-brighter (RWB) trends from color-magnitude diagrams, we classified sources based on values, considering indicative of a BWB trend and indicative of an RWB trend. For sources with values of between and 0.5, we assumed that there is no significant correlation (termed as NOT) between color and magnitude. The number of sources that exhibit trends of RWB, BWB, or NOT, along with their fractions relative to the sample size, is summarized in Table 3. The slope, intercept, and color trends of the color-magnitude diagrams g-r vs. r, r-i vs. i, and W1-W2 vs. W2 for the individual source are tabulated in Table 4.
For optical wavelength | For the mid-infrared wavelength | ||||||||||
Seyfert galaxies | RWB (%) | BWB (%) | NOT (%) | Seyfert galaxies | RWB (%) | BWB (%) | NOT (%) | Seyfert galaxies | RWB (%) | BWB (%) | NOT (%) |
[no. of sources] | for (g-r) color vs. magnitude | [no. of sources] | for (r-i) color vs. magnitude | [no. of sources] | for (W1-W2) color vs. magnitude | ||||||
gNLS1 [17] | 6 (35.29%) | 1 (5.88%) | 10 (58.82%) | gNLS1 [10] | 5 (50.00%) | 0 (0.00%) | 5 (50.00%) | gNLS1 [22] | 15 (68.18%) | 0 (00.00%) | 7 (31.82%) |
ngNLS1 [128] | 78 (60.94%) | 0 (0.00%) | 50 (39.06%) | ngNLS1 [74] | 55 (74.32%) | 0 (0.00%) | 19 (25.68%) | ngNLS1 [176] | 169 (96.02%) | 0 (00.00%) | 7 (3.98%) |
gBLS1 [10] | 8 (80.00%) | 0 (0.00%) | 2 (20.00%) | gBLS1 [6] | 3 (50.00%) | 0 (0.00%) | 3 (50.00%) | gBLS1 [10] | 3 (30.00%) | 1 (10.00%) | 6 (60.00%) |
The number of sources used to estimate the statistics are tabulated inside the bracket ‘[]’. |
4 Results and discussion
In this study, we systematically investigated flux and color variability for redshift-matched samples of gNLS1s, ngNLS1s, and gBLS1s in optical and mid-infrared wavelengths. For this purpose, we utilize high-cadence optical light curves in the g, r, and i bands from ZTF and mid-infrared light curves in the W1 and W2 bands from WISE. To analyze flux variability in the optical wavelength, we initially applied -test using Eq. 1 to the g, r, and i band light curves to determine the presence or absence of intrinsic variability. Subsequently, was calculated exclusively for the variable sources using equation 2, whereas F was calculated for the entire sample, including both variable and non-variable targets, using equation 3. The mean and median values of and are tabulated in Table 1.
From Table 1 and Fig. 2, it is evident that gBLS1s exhibit greater variability, as indicated by both and , compared to gNLS1s across the g, r, and i bands, whether considering mean or median values. This finding is consistent with previous studies on smaller and larger samples of gNLS1s and gBLS1s (Klimek et al., 2004; Ai et al., 2010, 2013; Rakshit & Stalin, 2017). As a class, NLS1s demonstrate lower optical variability amplitudes than BLS1s. Such a lower variability in NLS1s can be attributed to their higher Eddington ratios (REdd). Since NLS1s accrete at rates significantly higher than BLS1s (Boroson & Green, 1992; Peterson et al., 2000; Ojha et al., 2020), their optical emission is expected to have a relatively stronger thermal component from the accretion disk compared to the synchrotron emission (e.g., see Zhou et al., 2007; Paliya et al., 2014). However, the optical emission from AGNs, dominated by thermal processes, is less variable than the Doppler-boosted synchrotron emission from jets, as a result, the higher contribution of thermal emission in NLS1 due to its higher accretion rate can effectively suppress the amplitude of their optical variability (e.g., see Ojha et al., 2019). However, in contrast to gBLS1s and gNLS1s, ngNLS1s are scarcely variable across the g, r, and i bands with a mean DC () of 5% (see Table 1, see, e.g., Fig. 2), suggest their variability due to instabilities in the accretion disc. Furthermore, the stark contrast in variability characteristics between the gNLS1 and ngNLS1 samples reflected in their mean duty cycles of approximately 72% and 5%, respectively, can be attributed to the presence of relativistic jets in gNLS1s. These jets, which are responsible for the observed gamma-ray emission in gNLS1s, likely contribute significantly to the enhanced and more frequent variability exhibited by this subclass. The relativistic beaming effect associated with the jets amplifies both the emission and its variability, resulting in a markedly higher detection rate of flux variations compared to jet-absent ngNLS1s.
For the optical wavelength | For the mid-infrared wavelength | ||||||||||||
(g-r) color vs. magnitude | (r-i) color vs. magnitude | (W1-W2) color vs. magnitude | |||||||||||
SDSS Name | Galaxy | Slope | Intercept | Color | Slope | Intercept | Color | Slope | Intercept | Color | |||
type | trends | trends | trends | ||||||||||
J112758.87362028.4 | gNLS1 | 0.230.10 | 4.821.88 | 0.51 | RWB | 0.200.20 | 04.263.77 | 0.56 | RWB | 0.230.04 | 04.130.49 | 0.50 | RWB |
J130522.74511640.3 | gNLS1 | 0.520.05 | 9.040.87 | 0.55 | RWB | 0.400.10 | 06.771.65 | 0.54 | RWB | 1.090.07 | 14.350.80 | 0.65 | RWB |
J144318.56472556.7 | gNLS1 | 0.300.07 | 5.441.34 | 0.27 | NOT | 0.530.17 | 09.563.00 | 0.37 | NOT | 1.160.05 | 15.400.61 | 0.75 | RWB |
J000545.32063945.4 | ngNLS1 | 0.360.15 | 6.892.86 | 0.39 | NOT | 0.570.21 | 10.743.90 | 0.62 | RWB | 0.990.06 | 14.650.86 | 0.69 | RWB |
J000935.91295425.8 | ngNLS1 | 0.390.11 | 8.052.17 | 0.53 | RWB | 1.290.49 | 25.959.72 | 0.52 | RWB | 1.190.05 | 18.110.73 | 0.86 | RWB |
J002151.43322520.0 | ngNLS1 | 0.110.12 | 2.472.21 | 0.15 | NOT | 0.690.24 | 13.054.60 | 0.75 | RWB | 1.330.20 | 18.602.73 | 0.72 | RWB |
J110249.85525012.7 | gBLS1 | 0.310.05 | 6.270.91 | 0.56 | RWB | 0.260.12 | 05.402.27 | 0.68 | RWB | 0.450.03 | 07.230.49 | 0.63 | RWB |
J120951.76181006.9 | gBLS1 | 0.270.09 | 5.331.65 | 0.54 | RWB | 0.090.09 | 01.721.70 | 0.53 | RWB | 0.020.01 | 01.270.16 | 0.31 | NOT |
J164258.81394837.0 | gBLS1 | 0.170.01 | 03.230.18 | 0.79 | RWB | 0.030.02 | 0.150.38 | 0.20 | NOT | 0.270.06 | 04.070.68 | 0.51 | RWB |
Notes: A portion of this table is presented here to display its form and content, however, the entire table is available in the online electronic version. |
In contrast to the accretion disc based variability in ngNLS1s, the origin of resulted flux variabilities of gNLS1s and gBLS1s could either be due to accretion disc instabilities (e.g., Wiita et al., 1991; Chakrabarti & Wiita, 1993; Mangalam & Wiita, 1993) or fluctuations emerging within the jet (e.g., Marscher & Gear, 1985; Wagner & Witzel, 1995; Marscher, 2014). Therefore, we have adopted the variability amplitude cut of 0.4 mags suggested by Bauer et al. (2009) to distinguish between jet-based variability and accretion disc instability-based variability. Thus, we first checked for the variable targets in the samples of gNLS1s and gBLS1s, and found that 15 out of 17 variable gNLS1s (88%) and 10 out of 10 variable gBLS1s (100%) exhibit 0.4 mags for the r-band light-curves. For the g-/ i-band light-curves, such statistics are (89/75%) and (100/100%) for the samples of gNLS1s and gBLS1s, respectively. From this analysis, it appears that gBLS1s tend to be more variable compared to gNLS1s, which could be due to the larger observed contribution of jets more closely aligned towards the observer line of sight in gBLS1s (e.g. Safna et al., 2020). A comparatively small percentage of the variable targets with 0.4 mags in the sample of gNLS1s could also be due to a difference in their accretion rates as discussed above (see 2nd para). On the other hand, a similar kind of variability nature resulted from the samples of gNLS1s, ngNLS1s and gBLS1s in the mid-infrared W1 and W2 bands (see Table 1, Fig. 3), supporting a reprocessing of optical/UV emission from the central engine of Seyfert galaxies by the dusty torus to MIR wavelength. Furthermore, to strengthen the reprocessing scenario found here, we have checked the correlation of with and for optically variable targets having 0.1 in the mid-infrared wavelength for the samples of gNLS1s, ngNLS1s, and gBLS1s. This has resulted in a very good correlations ( 0.5) of with and for the r-band light curves in the samples of gNLS1s and gBLS1s both for W1 and W2 bands (see Fig. 4). However, such correlations resulted in a less significant equal to 0.46 and 0.43 for W1 and W2 bands for the correlation between and only in the sample of gNLS1s for the g-band light curves. Overall, for the g-band light curves, we have found stronger with mean for W1 and W2 bands () equal to 0.71 & 0.73 for the correlation of with and , respectively in gBLS1s compared to value equal to 0.45 & 0.63 of gNLS1s. Such statistics of found to be 0.70 & 0.87 for gBLS1s and 0.56 & 0.57 for gNLS1s when considering r-band light curves. Thus, it appears that values are consistently found to be stronger in the sample of gBLS1s as compared to the sample of gNLS1s for the correlation of with and . A comparatively less significant found in the sample of gNLS1s could be due to the higher accretion rate of gNLS1s which effectively suppresses the amplitude of variability. Contrary to the significant correlations found for the g-band and r-band light curves, such correlations are found to be weaker for the i-band light curves. On the other hand, no such correlation has been found in the sample of ngNLS1s, except for i-band light curves. Since i-band data are very scarcely sampled compared to g and r-band data, therefore, it is difficult the draw any conclusion based on the i-band data. The lack of any significant correlation in the sample of ngNLS1s could be either due to a small sample size which is 8 & 6 sources only for W1 and W2 bands (see Fig. 4) or due to large contamination in the variability amplitude by the host galaxies.
The relatively weak variability (duty cycle 5%) and the lack of significant correlation between the optical and infrared variability amplitudes observed in ngNLS1s may be attributed to their characteristically weak yet rapid variability, manifesting over short (intranight) timescales. Such rapid fluctuations are not well captured by large-scale time-domain surveys such as ZTF and WISE, which typically operate with cadences of a day or longer. This interpretation aligns well with the findings of the Super-Eddington Accreting Massive Black Holes (SEAMBH) reverberation mapping (RM) project (e.g., see Du et al., 2018; Dunlop et al., 2003), which systematically monitored a sample dominated by ngNLS1s. The SEAMBH project demonstrated that H reverberation lags in these high accretion rate AGNs are significantly shorter by factors of 2-6 than those predicted by the canonical relation. These shortened lags reflect the compact nature of the broad-line region in SEAMBHs and underscore the importance of high-cadence observations for capturing their rapid variability. Thus, the weak variability signatures observed in the present study probably reflect a cadence-induced bias rather than an absence of intrinsic activity.
On the other hand, as discussed in Sect. 3.4, studying the color variability of AGNs can provide an effective way to disentangle between quasi-thermal emission from the accretion disc and non-thermal synchrotron emission from the jet, which contributes to the overall observed emission. Therefore, for our samples of gNLS1s, ngNLS1s and gBLS1s, we generated color-magnitude (g-r vs. r & r-i vs. i) diagrams in optical regime and W1-W2 vs. W2 in MIR regime using their quasi-simultaneous data observed in respective bands. In Fig. 5, long-term color-magnitude plots for (g-r) vs. r, (r-i) vs. i, & (W1-W2) vs. W2, respectively, are presented in top, middle, and bottom for representative examples of the gNLS1-J130522.74511640.2 (left), ngNLS1-J144019.82502201.7 (middle), and gBLS1-J110249.85525012.7 (right) from the current sample, illustrating RWB trend both in optical and MIR wavelengths, respectively. In these figures, the black dashed line represents the best fit, obtained using orthogonal distance regression, accounting for uncertainties in both color and magnitude measurements.
Furthermore, considering the criterion of and set to identify stronger BWB and RWB, respectively from the color-magnitude diagrams, we have found 35% (6/17) and 61% (78/128) of gNLS1s and ngNLS1s, respectively, exhibiting RWB trend in (g-r) vs. r color-magnitude diagram (see Table 3), however, only a gNLS1-J150506.48032630 6% (1/17) showed BWB (see Table 4). Such an RWB trend is found for 50% (5/10) and 74% (55/74) in samples of gNLS1s and ngNLS1s, respectively, when considering the color-magnitude diagram (r-i) vs. i (see Table 3). Similarly, in the sample of gBLS1s, 80% (8/10) and 50% (3/6) exhibit the RWB trend in the color-magnitude diagrams (g-r) vs. r and (r-i) vs. i, respectively (see Table 3). Note that 10 out of 17 gNLS1s, 50 out of 128 ngNLS1s, and 2 out of 10 gBLS1s didn’t show strong correlation between (g-r) color and r-band magnitude (see Table 3). However, for the correlation between the (r-i) color and i-band magnitude, these numbers are 5 out of 10 gNLS1s, 19 out of 74 ngNLS1s, and 3 out of 6 gBLS1s (see Table 3).
On the other hand, similar to the optical wavelength, 68% (15/22), 96% (169/176), and 30% (3/10) of gNLS1s, ngNLS1s, and gBLS1s, respectively, showed the RWB trend in color magnitude diagram of (W1-W2) vs. W2 (see Table 3). Contrary to the RWB trend exhibited by most sources, only a gBLS1s-J122454.45212246.3 showed a BWB trend (see Table 3). The remaining 32% (7/22), 4% (7/176), and 60% (6/10) of gNLS1s, ngNLS1s, and gBLS1s, respectively, didn’t exhibit a strong correlation between (W1-W2) color and W2-band magnitude (see Table 3).
The so-called BWB and RWB color trends from the AGN in the optical wavelength attribute to domination of non-thermal emission from the jet over thermal emission from the accretion disc, and vice-versa, respectively (e.g., see Malkan, 1983; Pian et al., 1999; Rani et al., 2010; Sakata et al., 2010; Ikejiri et al., 2011; Mao & Zhang, 2016; Negi et al., 2022). The BWB trend that predominantly appeared in the BL Lacs class of AGN can be explained with the shock-in-jet model (see Türler, 2011) where electrons are accelerated to higher energies at the front of the shock, and since high-energy electrons lose their energy faster compared to low-energy electrons by radiative cooling process, thus making the high-energy band (in the current work g-band compared to r-band and r-band compared to i-band) more variable (Kirk et al., 1998; Mastichiadis & Kirk, 2002). It could also be explained by an energy injection of fresh electrons due to internal shocks in the relativistic shell into the emitting regions, leading to an increase in the number of high energy electrons and eventually shifting to higher energy (e.g., Spada et al., 2001; Zhang et al., 2002; Fiorucci et al., 2004). Apart from the energy injection scenario, the BWB trend can also be attributed to variations in the beaming factor, , of the emitting region. Changes in not only result in an apparent increase in observed flux, following the relation , but also cause a shift in the observed frequency, (e.g., Villata et al., 2004; Papadakis et al., 2007; Larionov et al., 2010).
On the other hand, RWB trend that appeared in the Seyfert galaxies, can be explained by the domination of the disc’s fluctuations over the jet’s fluctuations. Since long-term optical variability in Seyfert galaxies is exhibited by a mix of disc and jet emissions, if the accretion disc’s fluctuations are primarily responsible for the variability but jet contributes more in blue/UV, then any brightening of the disc may increase the domination of redder thermal emission, leading to RWB trend (e.g., see Malkan, 1983; Pian et al., 1999; Sakata et al., 2010). The RWB trend observed in Seyfert galaxies and FSRQs can also be accommodated in the shock-in-jet model. When a shock propagates through a jet and interacts with regions of higher electron density, stronger magnetic fields, or turbulence, the emission from that region intensifies. Initially, higher-energy radiation emerges first, followed by a delayed peak at lower frequencies (e.g. Valtaoja et al., 2002). In the later stages of a flare, the flux enhancement in the redder bands may become more pronounced as the contribution from shorter wavelengths stabilizes or declines, leading to an overall RWB trend. This effect is particularly relevant in Seyfert galaxies, where the jet emission is weaker, and contributions from the disc and dust are more significant. In FSRQs, the strong thermal emission from the accretion disc, which peaks in UV, plays a crucial role. As the synchrotron emission from the jet increases, typically peaks at lower frequencies, it can shift the overall spectral energy distribution to redder colors, reinforcing the RWB trend (e.g., see Gu et al., 2006).
Thus, based on the long-term optical color variability observed in the current samples of gNLS1s, ngNLS1s, and gBLS1s, the dominant mechanism driving their long-term variability appears to be associated with processes within the accretion disc. This variability is likely induced by instabilities within the accretion disc. However, on shorter timescales, particularly on the intra-day timescale, the observed variability may be primarily attributed to jet-related processes.
On the other hand, based on the mid-infrared color-magnitude analysis (W1-W2) vs. W2, a large percentage 68%, 96%, and 30%, of gNLS1s, ngNLS1s, and gBLS1s, respectively, exhibit an RWB trend similar to that exhibited in optical wavelength (see Table 3). The observed RWB trend in the MIR wavelength in the current sample of Seyfert galaxies can be attributed to the dominance of a variable nuclear component over the constant host galaxy component, where the increased brightness of the AGN leads to a stronger contribution from the AGN dust torus to the observed MIR emission, resulting in a redder (W1-W2) color (e.g., see Stern et al., 2012; Yang et al., 2018). Additionally, a similar color trend observed in optical and mid-infrared wavelengths for most of the sources also reconfirms the reprocessing scenario. Thus, long-term MIR color variability analysis of gNLS1s, ngNLS1s, and gBLS1s reconfirms the conclusions drawn based on optical wavelength for the current sample.
5 Conclusions
We conducted a systematic investigation of flux and color variability in a redshift-matched sample of Seyfert galaxies, comprising 23 gNLS1s, 190 ngNLS1s, and 10 gBLS1s across optical and mid-infrared wavelengths. This analysis utilizes high-cadence photometric observations from the ZTF in the g, r, and i bands, along with mid-infrared measurements in the W1 and W2 bands from the WISE. To ensure the robustness of the spectral analysis, we impose a criterion of at least five quasi-simultaneous observations (within 30 minutes) in the consecutive bands for each source. Most sources have observational baselines that extend up to 2000 days, allowing a comprehensive study of both long-term flux and color variability in this sample. The primary conclusions drawn from our analysis are as follows:
(i) The highest variability is exhibit by gBLS1s in both optical (g, r, i bands) and mid-infrared (W1, W2 bands), followed by gNLS1s, while ngNLS1s show the least variability across both wavelengths. The higher optical and mid-infrared variability observed in gBLS1s compared to gNLS1s suggests a more closely aligned jet toward the observer line of sight in gBLS1s. This could also be due to comparatively lower accretion rates of gBLS1, which contaminate less effectively non-thermal Doppler-boosted synchrotron emission of gBLS1s as compared to gNLS1s. However, the least variability (DC of 5%) shown by ngNLS1s among different samples of Seyfert galaxies studied here suggests their long-term variability due to instabilities in the accretion
disc.
(ii) The variability amplitude cut-off suggests that variability in ngNLS1s is primarily driven by accretion disk instabilities, while in gBLS1s and gNLS1s, both accretion disk fluctuations and jet-related processes contribute.
(iii) A similar kind of variability nature resulted from the samples of gNLS1s, ngNLS1s, and gBLS1s in the mid-infrared W1 and W2 bands, supporting a reprocessing of optical/UV emission from the central engine of Seyfert galaxies by the dusty torus to the MIR wavelength. The reprocessing scenario was confirmed based on a significant correlation found between optical and mid-infrared variability amplitudes for gNLS1s and gBLS1s in the g and r bands, while this correlation is weaker in the i band and nearly absent in ngNLS1s, except for their i band data, which are very scarcely sampled. Thus, no hints of a reprocessing scenario have been found for ngNLS1s based on flux variability analysis, which could be due to a large contamination in the variability amplitude by the host galaxies.
(iv) Long-term color-magnitude analysis reveals that most gNLS1s, ngNLS1s, and gBLS1s exhibit an RWB trend in both optical and MIR wavelengths. Hence, it suggests that the dominant mechanism driving their long-term variability appears to be associated with processes within the accretion disc. This variability is likely induced by instabilities within the accretion disc. However, on shorter timescales, particularly on the intra-day timescale, the observed variability may be primarily attributed to jet-related processes. However, we could not find any hint of a reprocessing scenario for ngNLS1s based on the nearly absent correlation found between optical and mid-infrared variability amplitudes, but a similar RWB color trend found for most of the sources in both optical and mid-infrared wavelengths strongly supports the reprocessing scenario.
(v) Thus, long-term mid-infrared flux and color variability analysis of gNLS1s, ngNLS1s, and gBLS1s reconfirms the conclusion drawn based on the optical wavelength for the current sample.
(vi) The present analysis of gNLS1s, ngNLS1s, and gBLS1s highlights the significance of multi-wavelength studies in disentangling the contributions of various emission mechanisms in Seyfert galaxies. The findings underscore the necessity of a more systematic approach involving larger sample sizes and high-cadence, multi-band observations across different timescales to refine our understanding of the physical processes governing AGN variability in future studies.
6 Acknowledgment
We acknowledge the support of the National Key R&D Program of China (2022YFF0503401) and the National Science Foundation of China (12133001). The work of LCH was supported by the National Science Foundation of China (12233001), the National Key R&D Program of China (2022YFF0503401) and China Manned Space Program (CMS-CSST-2025-A09). This research utilizes observational data collected with the 48-inch Samuel Oschin Telescope, as part of the Zwicky Transient Facility (ZTF) initiative. The ZTF project is funded by the National Science Foundation through Grant No. AST-2034437 and is a collaborative effort involving institutions such as Caltech, IPAC, the Weizmann Institute of Science, the Oskar Klein Center at Stockholm University, the University of Maryland, DESY and Humboldt University, the TANGO Consortium in Taiwan, the University of Wisconsin–Milwaukee, Trinity College Dublin, Lawrence Livermore, National Laboratory, and IN2P3 in France. The operational management of ZTF is handled by Caltech Optical Observatories (COO), IPAC, and the University of Washington. This study also incorporates data from the Wide Field Infrared Survey Explorer (WISE), a collaborative mission between the University of California, Los Angeles, and NASA’s Jet Propulsion Laboratory at Caltech, supported by the National Aeronautics and Space Administration.
Data availability
The optical and mid-infrared data utilized in the current study are publicly available in ZTF DR23 (https://www.ztf.caltech.edu/ztf-public-releases.html) and the final release of the WISE data (https://irsa.ipac.caltech.edu/Missions/wise.html).